Circular No. 7842

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002bk IN ANONYMOUS GALAXY W. M. Wood-Vasey, G. Aldering, and P. Nugent, Lawrence Berkeley

National Laboratory, report the discovery of an apparent supernova (mag 18.5) on unfiltered CCD images taken with the Oschin 1.2-m Schmidt telescope at Palomar on Feb. 9.28 UT as part of a joint collaboration with E. F. Helin, S. Pravdo, M. Hicks, and K. Lawrence (Near-Earth Asteroid Tracking project, Jet Propulsion Laboratory). SN 2002bk is located at α = 9^{h}02^{m}26^{s}.67, δ = −9^{o}17′38′′.7 (equinox J2000.0), which is 3′′.0 west and 6′′.5 north of the nucleus of its apparent host galaxy. The new object was confirmed at mag 19.1 on subsequent images taken with the same telescope on Mar. 3.3, but images taken on Jan. 11.5 (limiting mag 19.5) showed nothing at this position.

SUNGRAZING COMETS Further to IAUC 7841, D. Biesecker and D. Hammer report their mea-

surements for five more comets found on images taken with the C2 coronagraph onboard SOHO. The 1998 objects were found by S. Hönig, and the 1999 ones by R. Kracht, on SOHO website images. The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below. C/1999 K16 and C/1999 L9 belong to the Meyer group (cf. IAUC 7832), and the others to the Kreutz group.

Comet  1998 UT  α_{2000}  δ_{2000}  MPEC
C/1998 H5 Apr. 20.173  1^{h}57.^{m}8  +10^{o}22′  2002-D42
C/1998 H6  20.255  1 57.6  +10 22  2002-D42
Comet  1999 UT  α_{2000}  δ_{2000}  MPEC
C/1999 K16 May 26.576  4^{h}13.^{m}7  +22^{o}26′  2002-E05
C/1999 K17  27.754  4 17.0  +19 39  2002-E05
C/1999 L9  June 9.368  5 09.7  +24 20  2002-E05

V838 MONOCEROTIS Visual magnitude estimates: Feb. 11.510 UT, 7.3 (A. Pearce, Nedlands,

W. Australia); 15.753, 7.7 (A. Baransky, Kiev, Ukraine); 21.948, 8.0 (A. Kammerer, Ettlingen, Germany); 27.510, 8.7 (Pearce); Mar. 1.979, 8.2 (R. Y. Shida, São Paulo, Brazil); 3.769, 8.0 (K. Hornoch, Lelekovice, Czech Republic); 5.496, 7.2 (Pearce).

2002 March 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green